Developing and Using a Predetermined Overhead Rate

Assume that the following predictions were made for last year for one of the plants of Milliken & Company:

Total manufacturing overhead for the year

$ 15,000,000

Total machine hours for the year

1,200,000


Actual results for February were as follows:

Manufacturing overhead

$ 1,238,500

Machine hours

98,500


(a) Determine the predetermined overhead rate per machine hour. (Enter answer using two decimal places.)
$Answer



(b) Using the predetermined overhead rate per machine hour, determine the manufacturing overhead applied to Work-in-Process during February.
$Answer



(c) As of February 1, actual overhead was underapplied by $35,000. Determine the cumulative amount of any overapplied or underapplied overhead at the end of February.
$Answer
Answer #1

a) Predetermine overhead rate = 15000000/1200000 = 12.50 per MH

b) Overhead applied = 98500*12.50 = $1231250

c) Actual overhead = 1238500

Applied overhead = 1231250

Under applied overhead for February = 1238500-1231250 = 7250

Cumulative amount of under applied overhead at the end of Feb = 35000+7250 = 42250